Suppose that in a perfectly competitive market, demand is given by Q=59.0-P and supply is given by Q=P-28.0. The government imposes a per-unit excise tax of $1 on the good. What is the dead-weight loss? No units, no rounding.
Answer #1

Answer

the equilibrium before tax is at Qd=Qs

59-P=P-28

2P=87

P=43.5

Q=59-43.5

Q=15.5

========

after-tax the supply curve is:

Q=P-28

P=Q+28

adding tax

P=Q+28+1=Q+29

P=Q+29

Q=P-29

now the equilibrium at new supply =Qd

P-29+59-P

2P=88

P=44=price consumer pay

Price producer receives =P-tax=44-1=43

Q=59-44=15

==========

deadweight loss =0.5*tax per unit * change in quantity

=0.5*1*(15.5-15)

=$0.25

the deadweight loss is 0.25

===============